Multiply. Simplify your answer and write as a mixed or whole number.
6 1/2 x 4 2/3 = ?

Answer by sudhanshu_kmr(1152)

6 1/2 x 4 2/3
= 13/2 * 14/3
= 91/3
= 30 ( 1/3)